OpenStudy (anonymous):

The longest side of a triangle is two times the shortest side. The medium side is 3 meters more than the shortest side. Perimeter is 27 meters. Find the dimensions

OpenStudy (anonymous):

let the longest side=x medium side=y shortest side=z x=2z y=z+3 x+y+z=27 now solve

OpenStudy (anonymous):

Let x, y and z be the longest, medium and shortest sides respectively: x = 2z y = z+3 z = z x+y+z = 27, so (2z)+(z+3)+(z) = 27, 4z+3 = 27, z = 8 The rest should follow.